### v2

@@ 4.3 Unbalanced Markets
We extend our results for smooth distributions to unbalanced markets satisfying $n\leq m\leq O(n)$.
@@
Suppose that $n\leq m\leq O(n)$ and $\mathbb{D}$ is $\beta$-smooth for some constant $\beta>0$. When servers are adversarially chosen and requests are independently drawn from $\mathbb{D}$, an algorithm exists with a competitive ratio of $O(\sqrt{n})$ for $d=1$, $O(\sqrt{\log n})$ for $d=2$, and $O(1)$ for $d\geq 3$.
@@
Let ${\mathcal{A}}$ be the algorithm as per Lemma 8. By Theorem 2, an algorithm ${\mathcal{A}}^{\prime}$ exists satisfying $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)$ for every set of servers $S$ with $n\leq|S|\leq O(n)$. Combining the lower bounds for $\mathrm{OPT}(S,n)$ given by Lemma 4 concludes the proof. ∎
5 Stochastic Setting: Euclidean Space with General Distributions
In this section, we assume the metric space to be $[0,1]^{d}$ with the Euclidean distance. We do not make any assumptions on the distribution and achieve nearly optimal regret bounds for all $d\geq 1$ in unbalanced markets with an arbitrary number of excess servers. In particular, our results recover the nearly optimal regret bounds by [CKKZ23], which is established in balanced markets with servers and requests following different distributions. In the weaker setting studied by [CKKZ23], the regret bound is optimal for $d\neq 2$ and optimal up to a logarithmic factor for $d=2$, implying that our regret bounds are also nearly optimal.
@@ 5 Stochastic Setting: Euclidean Space with General Distributions
Suppose that $n\leq m$ and $\mathbb{D}$ is an arbitrary distribution over $[0,1]^{d}$. An algorithm ${\mathcal{A}}^{\prime}$ exists that, for every set of servers $S$ with $|S|=m$, admits a regret of
$\mathrm{Regret}({\mathcal{A}}^{\prime};S,n)\leq\begin{cases}O(n^{-\frac{1}{2}})&d=1\\ O(n^{-\frac{1}{2}}\log n)&d=2\\ O(n^{-\frac{1}{d}})&d\geq 3\end{cases}\,.$
@@
For an arbitrary distribution $\mathbb{D}$, by applying the same arguments as in the proof of Lemma 8 except that the upper bound for $\mathrm{OPT}(n)$ is given by Lemma 1 instead of Lemma 2, we obtain an algorithm ${\mathcal{A}}$ such that
$\mathrm{cost}_{{\mathcal{A}}}(n)\leq\begin{cases}O(\sqrt{n})&d=1\\ O(\sqrt{n}\log n)&d=2\\ O(n^{1-\frac{1}{d}})&d\geq 3\end{cases}\,.$
By Theorem 2, an algorithm ${\mathcal{A}}^{\prime}$ exists satisfying $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)$ for every set of servers $S$ with $|S|\geq n$. As a result, the regret of ${\mathcal{A}}^{\prime}$ equals
$\mathrm{Regret}({\mathcal{A}}^{\prime};S,n)=\frac{1}{n}\left(\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)-\mathrm{OPT}(S,n)\right)\leq\frac{\mathrm{cost}_{{\mathcal{A}}}(n)}{n}\,,$
@@
We briefly explain the reason behind the contrast that in unbalanced markets with $d\leq 2$, our approach admits a nearly optimal regret (Theorem 7) with a suboptimal competitive ratio (Theorem 6). In fact, our regret guarantee indicates that $\mathrm{cost}_{{\mathcal{A}}}(S,n)-\mathrm{OPT}(S,n)$ is small, yet we have no control over the lower bound for $\mathrm{OPT}(S,n)$ even with smooth distributions, which prevents us from achieving a nearly optimal competitive ratio.
@@ 6 Prediction Setting
In this section, we adopt the same setting and notations as in Section 3.2. We will present an algorithm with desirable consistency and robustness guarantees, whose performance also smoothly depends on the prediction error. Recall that $\eta$ denotes the prediction error and $\mathrm{OPT}$ denotes the optimal offline cost, and our main result in this section is presented as follows.
@@
Assume that $m=n$. Given an algorithm ${\mathcal{A}}$ that does not use predictions and is $\alpha$-competitive for a particular metric, there exists an algorithm ${\mathcal{A}}^{\prime}$ that leverages predictions and is $O(\min\{\alpha,1+\frac{\eta(\alpha+1)}{\mathrm{OPT}}\})$-competitive for the same metric. Moreover, if ${\mathcal{A}}$ is deterministic, then ${\mathcal{A}}^{\prime}$ is also deterministic.
The proof of Theorem 8 is deferred to Section 6.1, and we now apply Theorem 8 to the existing algorithms in the literature. Recall that without predictions, there exist deterministic $(2n-1)$-competitive algorithms for general metrics [KP93, KMV94, Rag16], a deterministic $O(\log n)$-competitive algorithm for the line metric [Rag18], a randomized $O(\log^{2}n)$-competitive algorithm for general metrics [BBGN14], and a randomized $O(\log n)$-competitive algorithm for doubling metrics [GL12]. As a corollary, we derive the following results.
@@ 6.1 Proof of Theorem 8
Recall that we have given a non-robust algorithm with the desired consistency guarantee in Theorem 3, and now we show how to robustify it, which leads to the desired algorithm.
A powerful machinery for deterministically combining several algorithms into one in a black-box manner is proposed by [FRR94] for the $k$-server problem and is subsequently applied by [LV21, ACE+23b] under the context of online algorithm with predictions. We will use the same approach to combine our consistent but non-robust algorithm with the given algorithm ${\mathcal{A}}$ that does not use predictions, which is by definition robust but inconsistent. Consequently, robustness and consistency are simultaneously achieved.
We adapted the framework for combining several online algorithms generalized by [ACE+23b] to online metric matching, which is a direct consequence of combining [ACE+23b, Theorem 18] and [ACE+23b, Lemma 21].
@@
$\mathrm{cost}_{{\mathcal{A}}^{\mathrm{com}}}(S,R)\leq\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathrm{cost}_{{\mathcal{A}}_{i}}(S,R)\}\,.$
We generalize Lemma 9 to combine randomized online algorithms in the following lemma.
@@
$\mathrm{cost}_{{\mathcal{A}}^{\mathrm{com}}}(S,R)\leq\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathrm{cost}_{{\mathcal{A}}_{i}}(S,R)\}\,.$
@@
For each ${\mathcal{A}}_{i}$, recall that ${\mathcal{A}}_{i}$ is randomized, and we use ${\mathcal{A}}_{i}(S,R\mid B)$ to denote the cost incurred by ${\mathcal{A}}_{i}$ with servers $S$ and requests $R$ conditional on the random bits used by ${\mathcal{A}}_{i}$ being $B=b_{1}b_{2}\cdots$. To construct our desired algorithm ${\mathcal{A}}^{\mathrm{com}}$, we first independently sample the random bits $B=b_{1}b_{2}\cdots$ used by ${\mathcal{A}}_{1},\ldots,A_{k}$ and then apply the combination scheme of Lemma 9 to combine ${\mathcal{A}}_{1},\ldots,{\mathcal{A}}_{k}$. By the guarantee of Lemma 9, for all $S$ and $R$, it holds that
$\mathrm{cost}_{{\mathcal{A}}^{\mathrm{com}}}(S,R)$ $\leq\mathbb{E}_{B}\left[\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathrm{cost}_{{\mathcal{A}}_{i}}(S,R\mid B)\}\right]$
$\leq\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathbb{E}_{B}[\mathrm{cost}_{{\mathcal{A}}_{i}}(S,R\mid B)]\}$
$=\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathrm{cost}_{{\mathcal{A}}_{i}}(S,R)\}\,,$
@@
Now, we are ready to prove Theorem 8. We only prove the statement for deterministic algorithms by using Theorem 3 and Lemma 9, and the statement for randomized algorithms can be established analogously by applying Theorem 3 and Lemma 10. Given a deterministic $\alpha$-competitive algorithm ${\mathcal{A}}$ that does not use predictions for a particular metric, by Theorem 3, there exists a deterministic $(1+\frac{\eta(\alpha+1)}{\mathrm{OPT}})$-competitive algorithm ${\mathcal{A}}^{\prime}$ that leverages predictions for the same metric. By applying Lemma 9 with $\gamma=2$ to combine ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$, we obtain a deterministic algorithm ${\mathcal{A}}^{\mathrm{com}}$ incurring a cost of $O(\min\{\mathrm{cost}_{{\mathcal{A}}}(S,R),\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,R)\})$ for all $S$ and $R$ with $|S|=|R|$, which implies that ${\mathcal{A}}^{\mathrm{com}}$ is $O(\min\{\alpha,1+\frac{\eta(\alpha+1)}{\mathrm{OPT}}\})$-competitive by the competitive ratios of ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$, concluding the proof.
@@ 7 Adversarial Setting
In this section, we focus on a more general setting where both servers and requests are adversarially chosen. This setting has been extensively studied in the past [Rag18, BBGN14], whereas all known guarantees are established in balanced markets. We show that the competitive ratio guarantees in balanced markets of algorithms satisfying certain natural conditions can be carried over to unbalanced markets, and the resulting competitive ratios in unbalanced markets only depend on the number of requests. Combined with the algorithms of prior work, we obtain competitive ratio guarantees in unbalanced markets with adversarial servers and requests, which in turn leads to improved guarantees in unbalanced markets with adversarial servers and stochastic requests. In particular, we show that both the $O(\log^{3}n)$-competitive algorithm of [MNP06] for general metrics and the $O(\log n)$-competitive algorithm of [GL12] for doubling metrics admit the same competitive ratios in unbalanced markets. The latter result also improves the $O(\sqrt{n})$ competitive ratio for $d=1$ in unbalanced markets with servers adversarially chosen and requests drawn from a smooth distribution in Theorem 6.
Given the execution of an online algorithm ${\mathcal{A}}$ on an instance with initial servers $S$, we use $N_{{\mathcal{A}}}(r\mid S)$ to denote the set of available servers closest to the request $r$ upon its arrival. We say that an online algorithm ${\mathcal{A}}$ is greedy if it always (randomly) matches a new request to one of the closest available servers, i.e., it matches $r$ to a server in $N_{{\mathcal{A}}}(r\mid S)$. In particular, every greedy algorithm is completely determined by a (randomized) tie-breaking rule such that when facing multiple closest available servers upon the arrival of a new request, the algorithm applies the tie-breaking rule to (randomly) decide who to match with. We say that the tie-breaking rule of a greedy algorithm ${\mathcal{A}}$ is local if it only depends on the set of closest available servers $N_{{\mathcal{A}}}(r\mid S)$. Moreover, the local tie-breaking rule of a greedy algorithm ${\mathcal{A}}$ is called non-increasing if the probability of selecting each server in $N_{{\mathcal{A}}}(r\mid S)$ does not increase by adding additional servers into $N_{{\mathcal{A}}}(r\mid S)$, i.e., for all sets of servers $S^{\prime},S^{\prime\prime}$ with $S^{\prime}\subseteq S^{\prime\prime}$, for every $s\in S^{\prime}$,
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime}\right\}\geq\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}\right\}\,.$
We show that the competitive ratio guarantee for every greedy algorithm with a local and non-increasing tie-breaking rule in balanced markets can be carried over to unbalanced markets.
@@
For an arbitrary metric, if a greedy algorithm ${\mathcal{A}}$ with a local and non-increasing tie-breaking rule is $\alpha$-competitive in balanced markets with adversarial servers and requests, then ${\mathcal{A}}$ is $\alpha$-competitive in unbalanced markets with adversarial servers and requests.
@@
Proof of Theorem 9.
For two sets of points $S$ and $R$ in the metric with $|S|\geq\mid R|$, denote $\mathrm{cost}_{{\mathcal{A}}}(S,R)$ as the expected cost of ${\mathcal{A}}$ when the set of initial servers is $S$ and the set of requests is $R$. Then, we claim that for all sets of points $S,S^{\prime},R$ in the metric with $S^{\prime}\subseteq S$ and $|S^{\prime}|\geq|R|$, $\mathrm{cost}_{{\mathcal{A}}}(S,R)\leq\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)$.
Before proving our claim, we first show the desired competitive ratio guarantee of ${\mathcal{A}}$ in unbalanced markets by applying the claim. Let $S$ be the set of initial servers and $R$ be the set of requests with $|S|\geq|R|=n$, and let $S^{\prime}\subseteq S$ satisfying $|S^{\prime}|=n$ and $\mathrm{OPT}(S,R)=\mathrm{OPT}(S^{\prime},R)$, i.e., $S^{\prime}$ consists of all matched servers in $S$ in the minimum cost matching between $S$ and $R$ that matches all requests in $R$. Since ${\mathcal{A}}$ is $\alpha$-competitive in balanced markets,
$\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)\leq\alpha\cdot\mathrm{OPT}(S^{\prime},R)=\alpha\cdot\mathrm{OPT}(S,R)\,.$
Given that $S^{\prime}\subseteq S$ and $|S^{\prime}|\geq|R|$, by our claim,
$\mathrm{cost}_{{\mathcal{A}}}(S,R)\leq\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)\leq\alpha\cdot\mathrm{OPT}(S,R)\,,$
@@
It remains to prove our claim, which we establish by induction on $|R|$. When $|R|=0$, we have $\mathrm{cost}_{{\mathcal{A}}}(S,R)=\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)=0$, and the statement straightforwardly holds. Assume for induction that the statement holds for $|R|=k$, and we show that the statement also holds for $|R|=k+1$. Given $S,S^{\prime},R$ with $S^{\prime}\subseteq S$ and $|S^{\prime}|\geq|R|=k+1$, let $r\in R$ be the first arrived request. We use $M_{{\mathcal{A}}}(r\mid S)$ and $M_{{\mathcal{A}}}(r\mid S^{\prime})$ to denote the (random) server that ${\mathcal{A}}$ matches $r$ to when the set of initial servers being $S$ and $S^{\prime}$, respectively. Since ${\mathcal{A}}$ is greedy, $M_{{\mathcal{A}}}(r\mid S)\in N_{{\mathcal{A}}}(r\mid S)$ and $M_{{\mathcal{A}}}(r\mid S^{\prime})\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. The following lemma states that either $N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})=\emptyset$ or $N_{{\mathcal{A}}}(r\mid S^{\prime})\subseteq N_{{\mathcal{A}}}(r\mid S)$.

### v3

@@ 4.3 Unbalanced Markets
We extend our results for balanced markets to unbalanced markets in the following theorem.
@@
Suppose that $m=\kappa n$ and $\mathbb{D}$ is a $\beta$-smooth distribution over $0,1]^{d}$ for $\kappa\geq 1$ and $\beta\geq 1$. Then, in the semi-stochastic setting, there exists an algorithm with a competitive ratio of $O(\beta\kappa\sqrt{n})$ for $d=1$, $O(\sqrt{\beta\kappa\log n})$ for $d=2$, and $O((\beta\kappa)^{\frac{1}{d}})$ for $d\geq 3$.
@@
Let $\mathcal{A}$} be the algorithm as per Lemma 7. By Theorem 2, there exists an algorithm $\mathcal{A}$$^{\prime}$ such that $\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}}^{\prime};S,n)\leq\mathrm{OPT}_{\mathbb{D}}(S,n)+\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};n)$ for every set of servers $S$ with $|S|\geq n$. Combining the lower bounds for $\mathrm{OPT}_{\mathbb{D}}(S,n)$ given by Lemma 3 concludes the proof. ∎
5 Semi-Stochastic Setting: Euclidean Metric with General Distributions
In this section, we consider the semi-stochastic setting and assume the metric space to be $0,1]^{d}$ with the Euclidean distance. For general distributions, we give an algorithm for unbalanced markets with an arbitrary number of excess servers, whose regret nearly matches the lower bounds established by [CKKZ23] for balanced markets and the stochastic setting.
@@ 5 Semi-Stochastic Setting: Euclidean Metric with General Distributions
Suppose that $\mathbb{D}$ is an arbitrary distribution over $0,1]^{d}$. Then, there exists an algorithm $\mathcal{A}$$^{\prime}$ with regret
$\mathrm{Regret}_{\mathbb{D}}({\mathcal{A}}^{\prime};S,n)\leq\begin{cases}O(n^{-\frac{1}{2}})&d=1\\ O(n^{-\frac{1}{2}}\sqrt{\log n})&d=2\\ O(n^{-\frac{1}{d}})&d\geq 3\end{cases}$
for every set of servers $S$ with $|S|\geq n$.
@@
By Lemma 7, there exists an algorithm $\mathcal{A}$} such that
$\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};n)\leq\begin{cases}O(\sqrt{n})&d=1\\ O(\sqrt{n\log n})&d=2\\ O(n^{1-\frac{1}{d}})&d\geq 3\end{cases}\,.$
By Theorem 2, we obtain an algorithm $\mathcal{A}$$^{\prime}$ such that $\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}}^{\prime};S,n)\leq\mathrm{OPT}_{\mathbb{D}}(S,n)+\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};n)$ for every set of servers $S$ with $|S|\geq n$. As a result, the regret of $\mathcal{A}$$^{\prime}$ equals
$\mathrm{Regret}_{\mathbb{D}}({\mathcal{A}}^{\prime};S,n)=\frac{1}{n}\left(\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}}^{\prime};S,n)-\mathrm{OPT}_{\mathbb{D}}(S,n)\right)\leq\frac{\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};n)}{n}\,$,
@@
We briefly explain the contrast that in unbalanced markets with $d\leq 2$, our algorithm exhibits nearly optimal regret (Theorem 7) with a suboptimal competitive ratio (Theorem 6). Intuitively, our regret analysis indicates that $\mathrm{cost}_{{\mathbb{D}}}({\mathcal{A}};S,n)-\mathrm{OPT}_{\mathbb{D}}(S,n)$ is small, yet we have no control over the magnitude of $\mathrm{OPT}_{\mathbb{D}}(S,n)$ even when $\mathbb{D}$} is smooth, which prevents us from achieving a better competitive ratio.
@@ 6 Prediction Setting
In this section, we focus on balanced markets and the adversarial setting. We assume that the algorithm is equipped with a prediction of request locations in advance. More formally, the prediction is given in the form of a set of predicted request locations $R^{\prime}$ with $|R^{\prime}|=n$, which is provided before the arrival of any request, and the prediction error is defined as $\eta:=\mathrm{OPT}(R^{\prime},R)$. We reiterate that the prediction error is not known to the algorithm. Throughout, we slightly abuse notations and use $\mathrm{OPT}:=\mathrm{OPT}(S,R)$ to denote the optimal offline cost. For an algorithm $\mathcal{A}$} that leverages predictions, we use ${\mathcal{A}}(R^{\prime})$ to indicate that the prediction provided to $\mathcal{A}$} is $R^{\prime}$.
Following the literature on algorithm with predictions [ACE+23b, BFGT24], we define the consistency of an algorithm as its competitive ratio when the prediction is perfect, i.e., when $\eta=0$, and the robustness of an algorithm as its worst-case competitive ratio over all predictions. In other words, if an algorithm $\mathcal{A}$} is $\alpha_{1}$-consistent and $\alpha_{2}$-robust, then for all $S$ and $R$, $\mathrm{cost}({\mathcal{A}}(R^{\prime});S,R)\leq\alpha_{1}\cdot\mathrm{OPT}(S,R)$ when $\mathrm{OPT}(R^{\prime},R)=0$, and $\mathrm{cost}({\mathcal{A}}(R^{\prime});S,R)\leq\alpha_{2}\cdot\mathrm{OPT}(S,R)$ for every prediction $R^{\prime}$. We provide our main theorem in this section as follows, which gives a consistent and robust algorithm. Moreover, the performance of the algorithm degrades smoothly as the prediction error increases.
@@
Suppose that $m=n$ and consider the adversarial setting. Then, given an algorithm $\mathcal{A}$} that does not use predictions and is $\alpha$-competitive, there exists an algorithm $\mathcal{A}$$^{\prime}$ that leverages predictions and is $O(\min\{\alpha,1+\frac{\eta(\alpha+1)}{\mathrm{OPT}}\})$-competitive. In particular, $\mathcal{A}$$^{\prime}$ is $O(1)$-consistent and $O(\alpha)$-robust. Moreover, if $\mathcal{A}$} is deterministic, then $\mathcal{A}$$^{\prime}$ is also deterministic.
The proof of Theorem 8 is deferred to Section 6.1, and we now apply Theorem 8 to the existing algorithms in the literature. Recall that without predictions, there exist deterministic (2n1)(2n-1)-competitive algorithms for general metrics [KP93, KMV94, Rag16], a deterministic $O(\log n)$-competitive algorithm for the line metric [Rag18], a randomized $O(\log^{2}n)$-competitive algorithm for general metrics [BBGN14], and a randomized $O(\log n)$-competitive algorithm for doubling metrics [GL12]. We derive the following corollary.
@@ 6.1 Proof of Theorem 8
We describe our high-level proof strategy as follows. We first apply Theorem 1 to give a consistent algorithm whose performance is optimal with perfect predictions and smoothly depends on the prediction error. Then, we robustify this algorithm by combining it with the provided algorithm $\mathcal{A}$}, which is by nature robust yet inconsistent.
We start by presenting a consistent algorithm using Theorem 1.
Lemma 8.
Given an $\alpha$-competitive algorithm $\mathcal{A}$} that does not use predictions, there exists an algorithm $\mathcal{A}$$^{\prime}$ that leverages predictions and is $(1+\frac{\eta(\alpha+1)}{\mathrm{OPT}})$-competitive. In particular, $\mathcal{A}$$^{\prime}$ is $1$-consistent. Moreover, if $\mathcal{A}$} is deterministic, then $\mathcal{A}$$^{\prime}$ is also deterministic.
Proof.
We construct $\mathcal{A}$$^{\prime}$ as follows. Given as input a set of servers $S$, a tuple of requests $R$, and a prediction of request locations $R^{\prime}$ with $|S|=|R|=|R^{\prime}|$ and $\eta=\mathrm{OPT}(R,R^{\prime})$, let $\mathcal{A}$$^{\prime}$ be the algorithm as per Theorem 1 by setting $I=R^{\prime}$. It follows that,
$\mathrm{cost}({\mathcal{A}}^{\prime}(R^{\prime});S,R)$ $\leq\mathrm{OPT}(S,R^{\prime})+\mathrm{cost}({\mathcal{A}};R^{\prime},R)\leq\mathrm{OPT}(S,R)+\mathrm{OPT}(R,R^{\prime})+\mathrm{cost}({\mathcal{A}};R^{\prime},R)$
Next, we robustify the consistent algorithm given by Lemma 8. A powerful machinery for deterministically combining several algorithms into one in a black-box manner is proposed by [FRR94] for the $k$-server problem and is subsequently applied by [LV21, ACE+23b] under the context of online algorithm with predictions. We will use the same machinery to combine our consistent but non-robust algorithm with the given algorithm $\mathcal{A}$} that does not use predictions, which is by definition robust but inconsistent, leading to a robust and consistent algorithm.
We adapted the framework for combining several online algorithms in [ACE+23b] to online metric matching, which is a direct consequence of combining [ACE+23b, Theorem 18 and Lemma 21].
@@
$\mathrm{cost}({\mathcal{A}}^{\mathrm{com}};S,R)\leq\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathrm{cost}({\mathcal{A}}_{i};S,R)\}.$
We generalize Lemma 9 to combine randomized algorithms in the following lemma.
@@
$\mathrm{cost}({\mathcal{A}}^{\mathrm{com}};S,R)\leq\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathrm{cost}({\mathcal{A}}_{i};S,R)\}.$
@@
For each randomized algorithm ${\mathcal{A}}_{i}$, we use $\mathrm{cost}({\mathcal{A}}_{i};S,R\mid B)$ to denote the cost incurred by ${\mathcal{A}}_{i}$ with servers $S$ and requests $R$ conditional on the random bits used by ${\mathcal{A}}_{i}$ being $B=b_{1}b_{2}\cdots$. To construct the algorithm ${\mathcal{A}}^{\mathrm{com}}$, we first sample the random bits $B=b_{1}b_{2}\cdots$ used by ${\mathcal{A}}_{1},\ldots,A_{k}$. Conditional on $B$, ${\mathcal{A}}_{1},\ldots,{\mathcal{A}}_{k}$ become deterministic, and hence we can apply Lemma 9 to combine ${\mathcal{A}}_{1},\ldots,{\mathcal{A}}_{k}$. By the guarantee of Lemma 9, for all $S$ and $R$, it holds that
$\mathrm{cost}({\mathcal{A}}^{\mathrm{com}};S,R)$ $\leq\mathbb{E}_{B}\left[\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathrm{cost}({\mathcal{A}}_{i};S,R\mid B)\}\right]$
$\leq\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathbb{E}_{B}[\mathrm{cost}({\mathcal{A}}_{i};S,R\mid B)]\}$
$=\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathrm{cost}({\mathcal{A}}_{i};S,R)\}\,$,
@@
Now, we are ready to prove Theorem 8.
Proof of Theorem 8.
We only prove the statement for deterministic algorithms by using Lemmas 8 and 9, and the statement for randomized algorithms can be established analogously by applying Lemmas 8 and 10. Given a deterministic $\alpha$-competitive algorithm $\mathcal{A}$} that does not use predictions, by Lemma 8, there exists a deterministic $(1+\frac{\eta(\alpha+1)}{\mathrm{OPT}})$-competitive algorithm $\mathcal{A}$$^{\prime}$ that leverages the prediction $R^{\prime}$. By applying Lemma 9 with $\gamma=2$ to combine $\mathcal{A}$} and $\mathcal{A}$$^{\prime}$, we obtain a deterministic algorithm ${\mathcal{A}}^{\mathrm{com}}$ such that
$\mathrm{cost}({\mathcal{A}}^{\mathrm{com}};S,R)\leq O(\min\{\mathrm{cost}({\mathcal{A}};S,R),\mathrm{cost}({\mathcal{A}}^{\prime}(R^{\prime});S,R)\})$
for all $S$ and $R$ with $|S|=|R|=|R^{\prime}|$, which implies that ${\mathcal{A}}^{\mathrm{com}}$ is $O(\min\{\alpha,1+\frac{\eta(\alpha+1)}{\mathrm{OPT}}\})$-competitive by the competitive ratios of $\mathcal{A}$} and $\mathcal{A}$$^{\prime}$, concluding the proof. ∎
@@ 7 Adversarial Setting
In this section, we focus on the adversarial setting without predictions. This setting has been extensively studied in the past [BBGN14, Rag18], whereas all known guarantees are established in balanced markets. We show that the competitive ratios in balanced markets of greedy algorithms satisfying certain natural conditions can be carried over to unbalanced markets. Combined with the algorithms of prior work, we obtain the first competitive ratio guarantees for unbalanced markets and the adversarial setting, which in turn leads to improved guarantees for unbalanced markets and the semi-stochastic setting. In particular, we show that both the $O(\log^{3}n)$-competitive algorithm of [MNP06] for general metrics and the $O(\log n)$-competitive algorithm of [GL12] for doubling metrics retain their competitive ratios in unbalanced markets. The latter result also improves the $O(\sqrt{n})$ competitive ratio for $d=1$ in Theorem 6.
Given the execution of an algorithm $\mathcal{A}$} on an instance with initial servers $S$, we use $N_{{\mathcal{A}}}(r\mid S)$ to denote the set of available servers closest to the request $r$ upon its arrival. We say that an algorithm $\mathcal{A}$} is greedy if it always (randomly) matches a new request to one of the closest available servers, i.e., it matches $r$ to a server in $N_{{\mathcal{A}}}(r\mid S)$. In particular, every greedy algorithm is completely determined by a (randomized) tie-breaking rule which, when facing multiple closest available servers upon the arrival of a new request $r$, (randomly) decide which server in $N_{{\mathcal{A}}}(r\mid S)$ to match $r$ with. We say that the tie-breaking rule of a greedy algorithm $\mathcal{A}$} is local if it only depends on the set of closest available servers $N_{{\mathcal{A}}}(r\mid S)$. Moreover, the local tie-breaking rule of a greedy algorithm $\mathcal{A}$} is called non-increasing if the probability of selecting each server in $N_{{\mathcal{A}}}(r\mid S)$ does not increase after adding additional servers into $N_{{\mathcal{A}}}(r\mid S)$, i.e., for all sets of servers $S^{\prime},S^{\prime\prime}$ with $S^{\prime}\subseteq S^{\prime\prime}$, for every $s\in S^{\prime}$,
[…]
We show that the competitive ratio of every greedy algorithm with a local and non-increasing tie-breaking rule in balanced markets is retained in unbalanced markets.
@@
For the adversarial setting, if a greedy algorithm $\mathcal{A}$} with a local and non-increasing tie-breaking rule is $\alpha$-competitive in balanced markets, then $\mathcal{A}$} is $\alpha$-competitive in unbalanced markets.
@@
Proof of Theorem 9..
Let $\mathcal{A}$} be a greedy algorithm with a local and non-increasing tie-breaking rule. We claim that for all sets of servers $S,S^{\prime}$ and tuple of requests $R$ with $S^{\prime}\subseteq S$ and $|S^{\prime}|\geq|R|$, $\mathrm{cost}({\mathcal{A}};S,R)\leq\mathrm{cost}({\mathcal{A}};S^{\prime},R)$. Before proving this claim, we first apply it to show the desired competitive ratio of $\mathcal{A}$} in unbalanced markets. Let $S$ be the set of servers and $R$ be the tuple of requests with $|S|\geq|R|=n$, and let $S^{\prime}\subseteq S$ satisfying $|S^{\prime}|=n$ and $\mathrm{OPT}(S,R)=\mathrm{OPT}(S^{\prime},R)$, i.e., $S^{\prime}$ consists of all matched servers in $S$ in the min-cost matching between $S$ and $R$ that matches all requests in $R$. Since $\mathcal{A}$} is $\alpha$-competitive in balanced markets,
$\mathrm{cost}({\mathcal{A}};S^{\prime},R)\leq\alpha\cdot\mathrm{OPT}(S^{\prime},R)=\alpha\cdot\mathrm{OPT}(S,R)\,.$
Given that $S^{\prime}\subseteq S$ and $|S^{\prime}|\geq|R|$, by the claim,
$\mathrm{cost}({\mathcal{A}};S,R)\leq\mathrm{cost}({\mathcal{A}};S^{\prime},R)\leq\alpha\cdot\mathrm{OPT}(S,R)\,$,
@@
It remains to prove the claim, which we establish by induction on $|R|$. When $|R|=0$, we have $\mathrm{cost}({\mathcal{A}};S,R)=\mathrm{cost}({\mathcal{A}};S^{\prime},R)=0$, and the statement straightforwardly holds. Assume for induction that the statement holds for $|R|=k$ with $k\geq 0$, and we show that the statement also holds for $|R|=k+1$. Given $S,S^{\prime},R$ with $S^{\prime}\subseteq S$ and $|S^{\prime}|\geq|R|=k+1$, let $r\in R$ be the first arrived request. We use $M_{{\mathcal{A}}}(r\mid S)$ and $M_{{\mathcal{A}}}(r\mid S^{\prime})$ to denote the (random) server that $\mathcal{A}$} matches $r$ to when the set of initial servers is $S$ and $S^{\prime}$, respectively. Since $\mathcal{A}$} is greedy, $M_{{\mathcal{A}}}(r\mid S)\in N_{{\mathcal{A}}}(r\mid S)$ and $M_{{\mathcal{A}}}(r\mid S^{\prime})\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. The following lemma asserts that either $N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})=\emptyset$ or $N_{{\mathcal{A}}}(r\mid S^{\prime})\subseteq N_{{\mathcal{A}}}(r\mid S)$.